# ConceptPose: Training-Free Zero-Shot Object Pose Estimation using Concept Vectors

### v2

@@ Abstract
Object pose estimation is a fundamental task in computer vision and robotics, yet most methods require extensive, dataset-specific training. Concurrently, large-scale vision language models show remarkable zero-shot capabilities. In this work, we bridge these two worlds by introducing ConceptPose, a framework for object pose estimation that is both training-free and model-free. ConceptPose leverages a vision-language-model (VLM) to create open-vocabulary 3D concept maps, where each point is tagged with a concept vector derived from saliency maps. By establishing robust 3D-3D correspondences across concept maps, our approach allows precise estimation of 6DoF relative pose. Without any object or dataset-specific training, our approach achieves state-of-the-art results on common zero shot relative pose estimation benchmarks, significantly outperforming existing methods by over 62% in ADD(-S) score, including those that utilize extensive dataset-specific training. We will release our code upon acceptance.
@@ 1 Introduction
Making machines see and think in 3D requires precise object understanding. In the era of embodied AI, object pose estimation has therefore become a critical building block for agents interacting with the physical world. The task is to determine an object’s 6D pose from an image. It enables key capabilities across robotic manipulation [wen2021catgrasp, wang2019densefusion, 6dpose_robotic_survey], augmented reality [ar_6dpose, Su2019DeepMO], and autonomous navigation [autonom_driving_6dpose]. However, classical approaches have long been constrained by a critical bottleneck: they require extensive, object-specific training, often depending on precise 3D CAD models and large datasets of ground-truth poses [hinterstoisser2013, xiang2018, wang2019, hodan2018]. This reliance on pre-trained models and curated data fundamentally limits their ability to adapt to novel objects and environments.
@@
To overcome this rigidity, the community has increasingly focused on category-level [wang2019, housecat6d2023] and zero shot relative pose estimation [oryon_corsetti2024, any6d2024, one2any_liu2025, uapose2025, unopose2025], which aim to handle previously unseen objects. In parallel, a powerful new trend has emerged with the advent of Large-scale Vision Foundation Models (VFMs) [dino_caron2021, clip_radford2021, siglip2023]. Trained on web-scale data, these models demonstrate unprecedented general-purpose understanding of semantics, geometry, and context—enabling strong zero-shot performance on tasks such as segmentation, detection, and captioning [2023CLIPDINOiser, 2024FindingDino, lidecap]. Naturally, recent 6D pose estimation methods have begun to exploit foundation models. Works such as FoundationPose [foundationpose2024], Oryon [oryon_corsetti2024], and Horyon [horyon_corsetti2024] leverage VFM backbones (e.g., DINO [dino_caron2021]) to extract robust 2D features, which are then used to learn dense correspondences within a trainable pose estimation pipeline. While this strategy achieves state-of-the-art results, it still depends on a crucial training phase: the VFM serves as a frozen feature extractor, but an additional ”head” or correspondence network must be trained on top. This limits true generalization and makes adaptation to newer, more powerful VFMs cumbersome.
@@ 2 Related Works
Model-based Object Pose Estimation rely on textured CAD models and object-specific training [hinterstoisser2013, xiang2018, wang2019, hodan2018]. Instance-level approaches [zebrapose2022, matchu2024, sam6d2024, gigapose2024] learn to estimate poses for specific instances, achieving high accuracy on known objects but failing to generalize to unseen ones. Training-free variants have been proposed for instance-level estimation [foundpose2024, caraffa2025], but these still require CAD models at test time, limiting their practical applicability. Category-level methods [wang2019, housecat6d2023, gspose2024, gcepose2025] relax the instance-specific constraint by learning category level representations (e.g., NOCS) that generalize across instances within a category. These methods still require extensive training and are mostly constrained to predefined object categories with available training data. These fundamental limitations motivate model-free, object-agnostic approaches.
@@
Model-free Object-agnostic Pose Estimation has emerged as a more practical alternative, eliminating the need for CAD models entirely. Multi-view approaches build 3D object representations from reference images or videos. Gen6D [gen6d2022], OnePose [onepose2022], and OnePose++ [oneposepp2022] use structure-from-motion to reconstruct sparse 3D point clouds and match 2D-3D correspondences. FS6D [fs6d2022] employs dense RGB-D prototypes with transformers for few-shot pose estimation. FoundationPose [foundationpose2024] constructs neural object fields from posed RGB-D images, achieving strong pose estimation through differentiable rendering. However, these multi-view methods require multiple reference views for object onboarding, which can be impractical in dynamic environments.
@@
Single-view reference methods have since gained increasing attention. They estimate pose from a single reference image, making them suitable for zero shot or few-shot scenarios. Learning-based methods like NOPE [nope2024], One2Any [one2any_liu2025], UNOPose [unopose2025], and UA-Pose [uapose2025] train networks to predict pose from single references. H-/Oryon [oryon_corsetti2024, horyon_corsetti2024] combine DINO [dino_caron2021] features with text embeddings of only the object name, limiting its further potential. All these approaches require dataset-specific training and cannot easily adapt to new foundation models without retraining.
@@
This motivates training-free alternatives that leverage foundation models directly. Classical methods like SIFT [lowe2004] and ObjectMatch [gu2023] perform geometric matching without learning, but lack semantic understanding and struggle with texture-less objects and large viewpoint changes. Recent work has begun exploring training-free pose estimation using VLMs. POPE [pope2024] employs DINOv2 features for zero-shot pose estimation through direct feature matching between reference and query images, demonstrating strong performance on OnePose benchmarks given non-standard evaluation protocols. While Any6D [any6d2024] claims training-free, they leverage image-to-3D generation models to reconstruct 3D object models from single reference images, then depends on FoundationPose [foundationpose2024]—a neural network pretrained on synthetic pose data—for final pose estimation. ConceptPose instead leverages VLM explainability methods to generate dense concept vectors from saliency maps, enabling semantic correspondence without learned pose networks or 3D reconstruction.
@@
Vision Language and Foundation Models like self-supervised foundation models DINO [dino_caron2021] and DINOv2 [dinov2_oquab2024] have been adopted in recent pose estimation methods  [oryon_corsetti2024, horyon_corsetti2024, foundpose2024, pope2024] as feature extractors for their emergent semantic properties. However, vision-language models (VLMs)—despite their rich semantic understanding from large-scale pretraining [clip_radford2021, li2022, li2023, zhai2023, siglip2023]—remain underutilized for pose estimation beyond feature extraction.
@@
While language-guided segmentation methods [ding2023, barsellotti2025talking, ren2024, groundingdino2024] excel at object-level localization through vision-language alignment, they struggle with diverse concepts humans naturally use for pose understanding, like object attributes, affordances and geometric characteristics. VLM explainability methods bridge this gap: techniques like Grad-CAM [selvaraju2020], originally designed for model interpretation, provide fine-grained spatial attention for arbitrary text queries that binary masks can’t. ConceptPose exploits this capability, repurposing VLM explainability from diagnosis to concept localization. By spatially grounding language-driven concepts, we enable training-free concept-level correspondence for pose estimation without learned matching networks or 3D model, unlocking VLMs for precise spatial reasoning.
@@ 3.2 Architecture
Saliency Map Extraction. For each concept $l_{i}\in\mathcal{L}$, we generate a spatial saliency map using a pre-trained VLM. We employ GradCAM [selvaraju2020] on the vision encoder to compute gradients with respect to the text prompt $l_{i}$.”. We first crop the input RGB image to the object’s bounding box and resize to the VLM’s input dimension. GradCAM computes class activation maps from the vision transformer’s output by weighted summation of gradients and activations, producing saliency maps at the input resolution. We finally resize the saliency map to the object’s bounding box size and pad back to the original image dimensions, yielding a $(L,H,W)$ saliency tensor where each channel highlights regions semantically aligned with concept $l_{i}$. To accelerate inference, we cache text embeddings across all frames.
@@
Pose Estimation. We estimate $\mathbf{T}_{\text{rel}}=(R_{\text{rel}},\mathbf{t}_{\text{rel}})$ using RANSAC [fischler1981]-based robust estimation: for each iteration, we sample a minimal set of correspondences and apply Umeyama’s [umeyama1989] closed-form solution to compute a similarity transformation $(R,\mathbf{t})$, then count inliers whose transformed distances fall below a threshold. RANSAC selects the transformation with maximum inliers, effectively optimizing for geometric consistency despite correspondence outliers. Finally, we apply ICP refinement using geometric nearest neighbor matching to locally optimize the alignment. The resulting transformation directly represents the camera-to-camera motion from anchor to query frame.
@@ 4.1 Experiment Setup
All experiments are conducted on AMD Ryzen 7 5800X 8-core CPU, and NVIDIA GeForce RTX 4060 Ti (16G), with 48GB of RAM. Our experiments uses PyTorch 2.0+ with CUDA 12.6. We employ SigLIP2-giant-opt-patch16-384 as the VLM and Gemini 2.5 Pro [gemini2025] for automated concept generation. Notably we are using FP16 precision for all operations to reduce memory usage on consumer-level hardware. We probe SigLIP’s vision encoder’s post_layernorm layer with pytorch-gradcam’s implementation. Unless otherwise specified, we report all results using our default prompt with $L=20$ for concept label generation, in our experiements, concepts are generated once and they applied to all instances within the category. We use deterministic RANSAC with a fixed random seed (seed=42) across all experiments to ensure reproducibility. For global ourlier removal, we discarding points beyond $\mu+2.5\sigma$ from the point cloud center. We perform 100,000 RANSAC iterations with a threshold of 0.01m.
@@ 4.2 Datasets
NOCS REAL275 [wang2019] dataset is a real-world category-level benchmark consisting of 2,754 RGB-D test frames across 6 indoor tabletop scenes. It contains 18 object instances from 6 categories (bottle, bowl, camera, can, laptop, mug), with 3 instances per category.
Toyota-Light (TYOL) [bop_2018] is a real-world dataset featuring 21 everyday object classes in realistic tabletop scenarios. The test split contains 21 scenes with varying illumination, clutter, and occlusion conditions Objects include household items such as mugs, plates, remote controls, and magazines.
YCB-Video (YCB-V) [xiang2018] is a highly cluttered dataset containing 21 unique objects from the YCB object set. YCB-V is characterized by significant object occlusions, challenging lighting variations, and dense clutter.
LINEMOD (LM) [hinterstoisser2013] is a classic 6D pose dataset featuring 15 unique objects including household items and tools. It includes challenging symmetrical objects (e.g., eggbox, glue, bowl).
@@ 4.3 Metrics
BOP Average Recall. Following the BOP Challenge protocol [hodan2018], we compute Average Recall (AR) across three complementary metrics: VSD (Visible Surface Discrepancy) evaluates pose quality via depth rendering with occlusion handling, MSSD (Maximum Symmetry-Aware Surface Distance) measures worst-case 3D point displacement, and MSPD (Maximum Symmetry-Aware Projection Distance) captures 2D reprojection accuracy. For VSD and MSSD, recall is computed across 10 thresholds ranging from 5% to 50% of the object diameter. For MSPD, thresholds span 5% to 50% of the image dimension. Each metric’s AR is the mean recall across its threshold range, and the final BOP Score averages the three AR values. All metrics incorporate symmetry handling to avoid penalizing perceptually equivalent poses.
@@
SIFT [lowe2004] 21.6 38.8 16.5 32.4 13.9 19.3 10.8 18.7 15.7 27.3
ObjectMatch [gu2023] 13.4 26.0 5.4 9.8 3.7 6.0 11.1 12.2 8.4 13.5
Oryon [oryon_corsetti2024] 34.9 46.5 22.9 34.1 12.8 19.4 20.4 25.3 22.8 31.3
Horyon [horyon_corsetti2024] 51.6 57.9 25.1 33.0 22.6 28.6 27.6 34.4 31.7 38.5
Any6D [any6d2024] 53.5 51.0 32.2 43.3 (42.9) (47.2)
One2Any [one2any_liu2025] 41.0 54.9 34.6 42.0 (37.8) (48.5)
@@ 4.4 Zero-shot Relative Pose Estimation Performance
Following the evaluation protocol proposed by Oryon [oryon_corsetti2024], we sample 2000 anchor-query pairs per dataset to assess relative pose estimation performance. For REAL275 and TYOL, we employ Oryon’s publicly available fixed splits, which contain 2000 cross-scene only pairs of the same object instance observed from different viewpoints. For YCB-Video and LINEMOD, we generate 2000 pairs via random sampling with a fixed seed filtered to BOP test targets (900 frames for YCB-V, 3000 for LINEMOD). YCB-Video pairs are cross-scene as objects appear in multiple scenes, while LINEMOD pairs are same-scene since each object appears in only one scene. We publish our generated YCB-V and LINEMOD anchor-query pairs using Oryon’s format convention for reproducibility.
Figure 3 shows extracted saliency maps and qualitative pose estimation results across the NOCS, YCB-V, and TYOL datasets. We show examples representing both geometrically rich objects with distinct semantic parts and symmetrical objects that pose significant challenges for correspondence-based methods. Table 1 compares ConceptPose against recent methods across four datasets. Note that Any6D [any6d2024] and One2Any [one2any_liu2025] were not evaluated on YCB-Video and LINEMOD, so we omit them from comparison on these benchmarks. For YCB-Video and LINEMOD, different sampling may introduce minor variance, but the performance gaps make comparisons meaningful.
For the recall of ADD(-S), ConceptPose outperforms all prior work on REAL275 (71.5 vs. 53.5), TYOL (55.0 vs. 32.2), YCB-Video (41.2 vs. 22.6), and LINEMOD (38.6 vs. 27.6), achieving 62.8% improvements on average over the strongest baselines. For BOP AR, ConceptPose also supasses all baselines on REAL275 (60.4 vs. 57.9), TYOL (51.6 vs. 43.3), YCB-Video (32.8 vs. 28.6), achieving a 14.3% improvement on average over the second best. It falls short slightly on LINEMOD (31.0 vs. 34.4 of Horyon [horyon_corsetti2024]) mainly due to the heavy occlusion in that dataset. Since we do not deploy any additional pipeline for occlusion handling whilst still being able to handle occlusion without training to a high degree, we believe this further validates the robustness of our approach. It is a good demonstration that concepts vectors can provide robust correspondence cues across diverse real-world scenarios.

### v3

@@ Abstract
Object pose estimation is a fundamental task in computer vision and robotics, yet most methods require extensive, dataset-specific training. Concurrently, large-scale vision language models show remarkable zero-shot capabilities. In this work, we bridge these two worlds by introducing ConceptPose, a framework for object pose estimation that is both training-free and model-free. ConceptPose leverages a vision-language-model (VLM) to create open-vocabulary 3D concept maps, where each point is tagged with a concept vector derived from saliency maps. By establishing robust 3D-3D correspondences across concept maps, our approach allows precise estimation of 6DoF relative pose. Without any object or dataset-specific training, our approach achieves state-of-the-art results on common zero shot relative pose estimation benchmarks, outperforming the strongest baseline by a relative 62% in average ADD(-S) score, including methods that utilize extensive dataset-specific training.
footnotetext: Project page: https://stevenlk.xyz/conceptpose.
@@ 1 Introduction
Making machines see and think in 3D requires precise object understanding. In the era of embodied AI, object pose estimation has therefore become a critical building block for agents interacting with the physical world. The task is to determine an object’s 6D pose from an image. It enables key capabilities across robotic manipulation [52, 50, 13], augmented reality [10, 44], and autonomous navigation [1]. However, classical approaches have long been constrained by a critical bottleneck: they require extensive, object-specific training, often depending on precise 3D CAD models and large datasets of ground-truth poses [17, 55, 51, 19]. This reliance on pre-trained models and curated data fundamentally limits their ability to adapt to novel objects and environments.
@@
To overcome this rigidity, the community has increasingly focused on category-level [51, 22] and zero shot relative pose estimation [8, 23, 30, 26, 32], which aim to handle previously unseen objects. In parallel, a powerful new trend has emerged with the advent of Large-scale Vision Foundation Models (VFMs) [5, 39, 48]. Trained on web-scale data, these models demonstrate unprecedented general-purpose understanding of semantics, geometry, and context—enabling strong zero-shot performance on tasks such as segmentation, detection, and captioning [54, 43, 27]. Naturally, recent 6D pose estimation methods have begun to exploit foundation models. Works such as FoundationPose [53], Oryon [8], and Horyon [7] leverage VFM backbones (e.g., DINO [5]) to extract robust 2D features, which are then used to learn dense correspondences within a trainable pose estimation pipeline. While this strategy achieves state-of-the-art results, it still depends on a crucial training phase: the VFM serves as a frozen feature extractor, but an additional ”head” or correspondence network must be trained on top. This limits true generalization and makes adaptation to newer, more powerful VFMs cumbersome.
@@ 2 Related Works
Model-based Object Pose Estimation rely on textured CAD models and object-specific training [17, 55, 51, 19]. Instance-level approaches [45, 20, 29, 36] learn to estimate poses for specific instances, achieving high accuracy on known objects but failing to generalize to unseen ones. Training-free variants have been proposed for instance-level estimation [38, 4], but these still require CAD models at test time, limiting their practical applicability. Category-level methods [51, 22, 3, 28, 47] relax the instance-specific constraint by learning category level representations (e.g., NOCS) that generalize across instances within a category. These methods still require extensive training and are mostly constrained to predefined object categories with available training data. These fundamental limitations motivate model-free, object-agnostic approaches.
@@
Model-free Object-agnostic Pose Estimation has emerged as a more practical alternative, eliminating the need for CAD models entirely. Multi-view approaches build 3D object representations from reference images or videos. Gen6D [33], OnePose [46], and OnePose++ [15] use structure-from-motion to reconstruct sparse 3D point clouds and match 2D-3D correspondences. FS6D [16] employs dense RGB-D prototypes with transformers for few-shot pose estimation. FoundationPose [53] constructs neural object fields from posed RGB-D images, achieving strong pose estimation through differentiable rendering. However, these multi-view methods require multiple reference views for object onboarding, which can be impractical in dynamic environments.
@@
Single-view reference methods have since gained increasing attention. They estimate pose from a single reference image, making them suitable for zero shot or few-shot scenarios. Learning-based methods like NOPE [35], One2Any [30], UNOPose [32], and UA-Pose [26] train networks to predict pose from single references. H-/Oryon [8, 7] combine DINO [5] features with text embeddings of only the object name, limiting its further potential. All these approaches require dataset-specific training and cannot easily adapt to new foundation models without retraining.
@@
This motivates training-free alternatives that leverage foundation models directly. Classical methods like SIFT [34] and ObjectMatch [14] perform geometric matching without learning, but lack semantic understanding and struggle with texture-less objects and large viewpoint changes. Recent work has begun exploring training-free pose estimation using VLMs. POPE [11] employs DINOv2 features for zero-shot pose estimation through direct feature matching between reference and query images, demonstrating strong performance on OnePose benchmarks given non-standard evaluation protocols. While Any6D [23] claims training-free, they leverage image-to-3D generation models to reconstruct 3D object models from single reference images, then depends on FoundationPose [53]—a neural network pretrained on synthetic pose data—for final pose estimation. ConceptPose instead leverages VLM explainability methods to generate dense concept vectors from saliency maps, enabling semantic correspondence without learned pose networks or 3D reconstruction.
@@
Vision Language and Foundation Models like self-supervised foundation models DINO [5] and DINOv2 [37] have been adopted in recent pose estimation methods  [8, 7, 38, 11] as feature extractors for their emergent semantic properties. However, vision-language models (VLMs)—despite their rich semantic understanding from large-scale pretraining [39, 25, 24, 56, 48]—remain underutilized for pose estimation beyond feature extraction.
@@
While language-guided segmentation methods [9, 2, 40, 31] excel at object-level localization through vision-language alignment, they struggle with diverse concepts humans naturally use for pose understanding, like object attributes, affordances and geometric characteristics. VLM explainability methods bridge this gap: techniques like Grad-CAM [41], originally designed for model interpretation, provide fine-grained spatial attention for arbitrary text queries that binary masks can’t. ConceptPose exploits this capability, repurposing VLM explainability from diagnosis to concept localization. By spatially grounding language-driven concepts, we enable training-free concept-level correspondence for pose estimation without learned matching networks or 3D model, unlocking VLMs for precise spatial reasoning.
@@ 3.2 Architecture
Saliency Map Extraction. For each concept $l_{i}\in\mathcal{L}$, we generate a spatial saliency map using a pre-trained VLM. We employ GradCAM [41] on the vision encoder to compute gradients with respect to the text prompt $l_{i}$.”. We first crop the input RGB image to the object’s bounding box and resize to the VLM’s input dimension. GradCAM computes class activation maps from the vision transformer’s output by weighted summation of gradients and activations, producing saliency maps at the input resolution. We finally resize the saliency map to the object’s bounding box size and pad back to the original image dimensions, yielding a $(L,H,W)$ saliency tensor where each channel highlights regions semantically aligned with concept $l_{i}$. To accelerate inference, we cache text embeddings across all frames.
@@
Pose Estimation. We estimate $\mathbf{T}_{\text{rel}}=(R_{\text{rel}},\mathbf{t}_{\text{rel}})$ using RANSAC [12]-based robust estimation: for each iteration, we sample a minimal set of correspondences and apply Umeyama’s [49] closed-form solution to compute a similarity transformation $(R,\mathbf{t})$, then count inliers whose transformed distances fall below a threshold. RANSAC selects the transformation with maximum inliers, effectively optimizing for geometric consistency despite correspondence outliers. Finally, we apply ICP refinement using geometric nearest neighbor matching to locally optimize the alignment. The resulting transformation directly represents the camera-to-camera motion from anchor to query frame.
@@ 4.1 Experiment Setup
All experiments are conducted on AMD Ryzen 7 5800X 8-core CPU, and NVIDIA GeForce RTX 4060 Ti (16G), with 48GB of RAM. Our experiments uses PyTorch 2.0+ with CUDA 12.6. We employ SigLIP2-giant-opt-patch16-384 as the VLM and Gemini 2.5 Pro [6] for automated concept generation. Notably we are using FP16 precision for all operations to reduce memory usage on consumer-level hardware. We probe SigLIP’s vision encoder’s post_layernorm layer with pytorch-gradcam’s implementation. Unless otherwise specified, we report all results using our default prompt with $L=20$ for concept label generation, in our experiements, concepts are generated once and they applied to all instances within the category. We use deterministic RANSAC with a fixed random seed (seed=42) across all experiments to ensure reproducibility. For global ourlier removal, we discarding points beyond $\mu+2.5\sigma$ from the point cloud center. We perform 100,000 RANSAC iterations with a threshold of 0.01m.
@@ 4.2 Datasets
NOCS REAL275 [51] dataset is a real-world category-level benchmark consisting of 2,754 RGB-D test frames across 6 indoor tabletop scenes. It contains 18 object instances from 6 categories (bottle, bowl, camera, can, laptop, mug), with 3 instances per category.
Toyota-Light (TYOL) [18] is a real-world dataset featuring 21 everyday object classes in realistic tabletop scenarios. The test split contains 21 scenes with varying illumination, clutter, and occlusion conditions Objects include household items such as mugs, plates, remote controls, and magazines.
YCB-Video (YCB-V) [55] is a highly cluttered dataset containing 21 unique objects from the YCB object set. YCB-V is characterized by significant object occlusions, challenging lighting variations, and dense clutter.
LINEMOD (LM) [17] is a classic 6D pose dataset featuring 15 unique objects including household items and tools. It includes challenging symmetrical objects (e.g., eggbox, glue, bowl).
@@ 4.3 Metrics
BOP Average Recall. Following the BOP Challenge protocol [19], we compute Average Recall (AR) across three complementary metrics: VSD (Visible Surface Discrepancy) evaluates pose quality via depth rendering with occlusion handling, MSSD (Maximum Symmetry-Aware Surface Distance) measures worst-case 3D point displacement, and MSPD (Maximum Symmetry-Aware Projection Distance) captures 2D reprojection accuracy. For VSD and MSSD, recall is computed across 10 thresholds ranging from 5% to 50% of the object diameter. For MSPD, thresholds span 5% to 50% of the image dimension. Each metric’s AR is the mean recall across its threshold range, and the final BOP Score averages the three AR values. All metrics incorporate symmetry handling to avoid penalizing perceptually equivalent poses.
@@
SIFT [34] 21.6 38.8 16.5 32.4 13.9 19.3 10.8 18.7 15.7 27.3
ObjectMatch [14] 13.4 26.0 5.4 9.8 3.7 6.0 11.1 12.2 8.4 13.5
Oryon [8] 34.9 46.5 22.9 34.1 12.8 19.4 20.4 25.3 22.8 31.3
Horyon [7] 51.6 57.9 25.1 33.0 22.6 28.6 27.6 34.4 31.7 38.5
Any6D [23] 53.5 51.0 32.2 43.3 (42.9) (47.2)
One2Any [30] 41.0 54.9 34.6 42.0 (37.8) (48.5)
@@ 4.4 Zero-shot Relative Pose Estimation Performance
Following the evaluation protocol proposed by Oryon [8], we sample 2000 anchor-query pairs per dataset to assess relative pose estimation performance. For REAL275 and TYOL, we employ Oryon’s publicly available fixed splits, which contain 2000 cross-scene only pairs of the same object instance observed from different viewpoints. For YCB-Video and LINEMOD, we generate 2000 pairs via random sampling with a fixed seed filtered to BOP test targets (900 frames for YCB-V, 3000 for LINEMOD). YCB-Video pairs are cross-scene as objects appear in multiple scenes, while LINEMOD pairs are same-scene since each object appears in only one scene. We publish our generated YCB-V and LINEMOD anchor-query pairs using Oryon’s format convention for reproducibility.
Figure 3 shows extracted saliency maps and qualitative pose estimation results across the NOCS, YCB-V, and TYOL datasets. We show examples representing both geometrically rich objects with distinct semantic parts and symmetrical objects that pose significant challenges for correspondence-based methods. Table 1 compares ConceptPose against recent methods across four datasets. Note that Any6D [23] and One2Any [30] were not evaluated on YCB-Video and LINEMOD, so we omit them from comparison on these benchmarks. For YCB-Video and LINEMOD, different sampling may introduce minor variance, but the performance gaps make comparisons meaningful.
For the recall of ADD(-S), ConceptPose outperforms all prior work on REAL275 (71.5 vs. 53.5), TYOL (55.0 vs. 32.2), YCB-Video (41.2 vs. 22.6), and LINEMOD (38.6 vs. 27.6), achieving 62.8% improvements on average over the strongest baselines. For BOP AR, ConceptPose also supasses all baselines on REAL275 (60.4 vs. 57.9), TYOL (51.6 vs. 43.3), YCB-Video (32.8 vs. 28.6), achieving a 14.3% improvement on average over the second best. It falls short slightly on LINEMOD (31.0 vs. 34.4 of Horyon [7]) mainly due to the heavy occlusion in that dataset. Since we do not deploy any additional pipeline for occlusion handling whilst still being able to handle occlusion without training to a high degree, we believe this further validates the robustness of our approach. It is a good demonstration that concepts vectors can provide robust correspondence cues across diverse real-world scenarios.